# Multiscale scattered data analysis in samplet coordinates

### v1

@@ 6.1. Franke’s function in two spatial dimensions
on the unit square $\Omega=[0,1]^{2}$. To this end, we generate 10 sets of nested points within this domain, distributing them on a regular grid with size $h_{j}=2^{-j}$. The number of points per level and the corresponding fill-distances are provided in Table 1.
@@
We have conducted computations using a-posteriori compression thresholds of $\kappa=10^{-4},10^{-6},10^{-8}$. For $\kappa=10^{-4}$, the error remains larger than $10^{-6}$, due to an increased compression error. Setting $\kappa=10^{-6}$, we have achieved an error of the order of $10^{-8}$. Further reducing the threshold to $\kappa=10^{-8}$ results in an error of $10^{-9}$. Even so, the latter results in a significantly worse compression rate of the generalized Vandermonde matrices.
Therefore, we detail the results using $\kappa=10^{-6}$ for the threshold. We approximate the solution with the ${C}^{1}$-smooth Matérn-3/2 RBF and the ${C}^{2}$-smooth Matérn-5/2 RBF and different values of $\gamma$. The error is evaluated at a fine grid with mesh-size $h=2^{-11}$. The accuracy of the conjugate gradient method is set to $10^{-6}$. Increasing the accuracy does not result in smaller errors but increases the number of iterations. Interpolation error, order and compression rate (%nz) as well as (non-preconditioned) conjugate gradient iterations (CG) are presented in the Tables 25.
@@
4 $6.76\cdot 10^{-3}$ 2.23 47 10 $1.15\cdot 10^{-3}$ 3.00 47 34
5 $1.48\cdot 10^{-3}$ 2.19 22 10 $2.51\cdot 10^{-4}$ 2.53 22 39
6 $3.45\cdot 10^{-4}$ 2.10 7.8 10 $6.08\cdot 10^{-5}$ 2.05 7.6 39
7 $8.30\cdot 10^{-5}$ 2.06 2.3 10 $1.30\cdot 10^{-5}$ 2.22 2.3 39
8 $1.46\cdot 10^{-5}$ 2.51 0.62 10 $1.14\cdot 10^{-6}$ 3.51 0.62 39
9 $3.14\cdot 10^{-6}$ 2.22 0.16 10 $1.04\cdot 10^{-7}$ 3.45 0.16 39
10 $7.89\cdot 10^{-7}$ 1.99 0.04 10 $4.47\cdot 10^{-8}$ 1.21 0.04 39
Table 2. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
@@
2 $4.67\cdot 10^{-2}$ 3.37 36 17 $4.51\cdot 10^{-2}$ 3.43 36 20
3 $7.97\cdot 10^{-3}$ 2.55 43 41 $7.72\cdot 10^{-3}$ 2.55 43 48
4 $5.70\cdot 10^{-4}$ 3.81 47 64 $3.60\cdot 10^{-4}$ 4.43 47 97
5 $1.09\cdot 10^{-4}$ 2.38 23 90 $5.75\cdot 10^{-5}$ 3.31 25 156
6 $2.38\cdot 10^{-5}$ 2.19 8.2 96 $1.16\cdot 10^{-5}$ 2.31 8.9 187
7 $4.57\cdot 10^{-6}$ 2.38 2.5 98 $2.07\cdot 10^{-6}$ 2.82 2.7 195
8 $4.77\cdot 10^{-7}$ 3.26 0.68 99 $2.27\cdot 10^{-7}$ 3.19 0.74 197
9 $5.97\cdot 10^{-8}$ 3.00 0.18 99 $7.85\cdot 10^{-8}$ 1.53 0.19 199
10 $4.63\cdot 10^{-8}$ 0.37 0.05 99 $7.30\cdot 10^{-8}$ 0.10 0.05 200
Table 3. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
@@
2 $1.12\cdot 10^{-1}$ 2.28 36 8 $4.51\cdot 10^{-2}$ 1.39 36 16
3 $2.27\cdot 10^{-2}$ 2.14 43 9 $7.73\cdot 10^{-3}$ 2.54 43 32
4 $4.54\cdot 10^{-3}$ 2.32 47 10 $6.92\cdot 10^{-4}$ 3.48 47 42
5 $1.00\cdot 10^{-3}$ 2.18 22 10 $1.49\cdot 10^{-4}$ 2.88 22 48
6 $2.49\cdot 10^{-4}$ 2.01 7.6 10 $3.42\cdot 10^{-5}$ 2.12 7.7 48
7 $6.13\cdot 10^{-5}$ 2.02 2.3 10 $6.88\cdot 10^{-6}$ 2.32 2.3 48
8 $6.06\cdot 10^{-6}$ 3.34 0.62 9 $8.31\cdot 10^{-7}$ 3.05 0.62 48
9 $1.15\cdot 10^{-6}$ 2.40 0.16 9 $6.15\cdot 10^{-8}$ 3.76 0.16 48
10 $3.07\cdot 10^{-7}$ 1.91 0.04 9 $1.57\cdot 10^{-8}$ 1.97 0.04 48
Table 4. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
@@
2 $4.41\cdot 10^{-2}$ 3.44 36 20 $4.69\cdot 10^{-2}$ 4.29 36 25
3 $7.13\cdot 10^{-3}$ 2.63 43 57 $7.47\cdot 10^{-3}$ 2.65 43 82
4 $3.29\cdot 10^{-4}$ 4.43 47 97 $1.63\cdot 10^{-4}$ 2.19 47 172
5 $6.14\cdot 10^{-5}$ 2.42 23 139 $3.96\cdot 10^{-5}$ 2.04 24 298
6 $1.22\cdot 10^{-5}$ 2.33 8.0 151 $1.34\cdot 10^{-5}$ 1.56 8.8 370
7 $2.11\cdot 10^{-6}$ 2.53 2.4 157 $2.06\cdot 10^{-6}$ 2.70 2.7 402
8 $2.57\cdot 10^{-7}$ 3.03 0.70 159 $2.39\cdot 10^{-7}$ 3.10 0.73 416
9 $4.06\cdot 10^{-8}$ 2.66 0.17 161 $7.56\cdot 10^{-8}$ 1.66 0.19 430
10 $3.04\cdot 10^{-8}$ 0.42 0.04 162 $7.01\cdot 10^{-8}$ 0.11 0.05 430
Table 5. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
The tables clearly demonstrate the expected exponential decay of the approximation error with increasing level for both error measures and both RBFs under consideration. Furthermore, we see the effect of a larger value of $\gamma$, which to a greater lengthscale parameter of the RBF, resulting in a larger condition number. For increasing $\gamma$, we require successively more iterations for the conjugate gradient method to achieve the prescribed accuracy, with up to 430 iterations for $\gamma=1$ and the Matérn-5/2 RBF. The deterioration of the convergence for $\gamma=0.75,1$ on the last level is caused by the samplet compression error, which is a directly to the chosen compression threshold $\kappa$.
Figure 2 shows a visualization of the solution (left) and the residuals (right) at levels $\ell=1,4,7,10$ for $\gamma=1$ using the Matérn-3/2 RBF.
@@ 6.2. Non-smooth function
We consider the L-shaped domain $\Omega=\left[-\frac{1}{2},\frac{1}{2}\right]^{2}\setminus\left(0,\frac{1}{2}\right]^{2}$ where we introduce polar coordinates $x=r\cos\phi$, $y=r\sin\phi$, with $r\geq 0$ and $\phi\in[\pi/2,2\pi]$. In this test, we aim at interpolating the function
$u(r,\phi)=-r^{\frac{2}{3}}\sin\left(\frac{2\phi-\pi}{3}\right),$
which solves the Laplace equation $\Delta u=0$. We select $h_{j}=2^{-j}$ as before. Given the reduced smoothness of the function, we choose the ${C}^{0}$-smooth Matérn-1/2 RBF for its approximation. The error is evaluated at a fine grid with mesh-size $h=2^{-9}$ and the accuracy of the conjugate gradient method is set to $10^{-6}$.
We have conducted experiments using $\kappa=10^{-4},10^{-6},10^{-8}$. While the magnitude of the errors remained consistent across these values, the compression rate in the final level increased as $\kappa$ decreased. Specifically for $\kappa=10^{-4}$, the compression rate is around $0.13\%$, for $\kappa=10^{-6}$, it is about $0.30\%$ and for $\kappa=10^{-8}$ it is approximately $0.43\%$. Consequently, we chose a higher threshold to reduce computational cost and storage. The number of points per level and the corresponding fill-distances are provided in Table 6.
@@
1 $4.37\cdot 10^{-2}$ $5.98\cdot 10^{-2}$ 52 9
2 $1.38\cdot 10^{-2}$ $3.17\cdot 10^{-2}$ 1.67 0.92 49 21
3 $5.33\cdot 10^{-3}$ $2.04\cdot 10^{-2}$ 1.37 0.64 38 28
4 $1.83\cdot 10^{-3}$ $1.29\cdot 10^{-2}$ 1.54 0.66 18 34
5 $5.88\cdot 10^{-4}$ $8.15\cdot 10^{-3}$ 1.64 0.67 6.2 36
6 $1.85\cdot 10^{-4}$ $5.04\cdot 10^{-3}$ 1.67 0.69 1.8 37
7 $5.87\cdot 10^{-5}$ $3.11\cdot 10^{-3}$ 1.66 0.70 0.51 37
8 $1.89\cdot 10^{-5}$ $1.61\cdot 10^{-3}$ 1.63 0.95 0.13 37
@@
1 $3.39\cdot 10^{-2}$ $6.06\cdot 10^{-2}$ 48 13
2 $1.27\cdot 10^{-2}$ $3.94\cdot 10^{-2}$ 1.41 0.62 49 28
3 $4.15\cdot 10^{-3}$ $2.50\cdot 10^{-2}$ 1.62 0.66 37 49
4 $1.29\cdot 10^{-3}$ $1.58\cdot 10^{-2}$ 1.68 0.66 18 73
5 $3.97\cdot 10^{-4}$ $9.94\cdot 10^{-3}$ 1.70 0.67 6.0 93
6 $1.23\cdot 10^{-4}$ $6.13\cdot 10^{-3}$ 1.70 0.70 1.8 102
7 $3.83\cdot 10^{-5}$ $3.72\cdot 10^{-3}$ 1.68 0.72 0.12 104
8 $1.20\cdot 10^{-5}$ $1.95\cdot 10^{-3}$ 1.68 0.76 0.09 105
Table 7. Results for the interpolation of the solution of the Laplace equation on the L-shaped domain using the ${C}^{0}$-smooth Matérn-1/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-4}$ as samplet compression parameters.
The compression rates are similar to the previous example, also in case of the less smooth Matérn-1/2 RBF. The approximation error and the corresponding order are reduced, however. The number of conjugate gradient iterations stays rather low with a maximum of 37 iterations for $\gamma=0.5$ and of 105 iterations for $\gamma=1$. This can be explained by the slower decay of the Matérn-1/2 kernel’s eigenvalues, resulting in less ill-conditioned generalized Vandermonde matrices.
@@
Figure 3. Solution (left) and corresponding residuals (right) for the interpolation of the solution to the Laplace equation on the L-shaped domain for levels $\ell=2,4,6,8$ using $\gamma=1$ and $\kappa=10^{-4}$.
@@ 6.3. Function on a point cloud in three spatial dimensions
As final example, we consider a function defined on a uniform resampling of the Lucy surface model of the Stanford 3D Scanning Repository. To construct the local approximation spaces, we selected random subsamples of the resulting point cloud, ensuring a specific cardinality for each set. We start with an initial set of 20 points, and at each subsequent level, the cardinality is increased by a factor of 16. In this study, we present results for both nested and non-nested point sets. For the nested case, the indices of the points are randomly shuffled, and subsamples are obtained by selecting the first $n$ points, where $n$ represents the desired cardinality. In contrast, for the non-nested case, each subsample is randomly selected from the entire dataset independently, i.e., without enforcing the inclusion of points from previous levels. Unlike previous tests, this experiment is more general, as the points are not arranged on a uniform grid, they are not necessarily nested, and they are not necessarily quasi-uniform. As test function, we e interpolate
$f({\boldsymbol{x}})=\frac{1}{\sqrt[4]{\|{\boldsymbol{x}}-{\boldsymbol{x}}_{0}\|_{2}}+10^{-4}}$
The point ${\boldsymbol{x}}_{0}$ is selected near the torch that Lucy holds in her left hand. To optimize performance, we use a relatively small value for the parameter $\gamma$, reduce the number of vanishing moments, and set a higher a-posteriori threshold. These adjustments help decrease both computational time and the number of non-zero entries. However, a very small $\gamma$ can compromise the precision, while a very large $\gamma$ increases the number of conjugate gradient iterations, thereby increasing computational time. A good trade-off is given by $\gamma=0.2$. Furthermore, we employed a diagonally preconditioned conjugate gradient method with accuracy $10^{-6}$ to solve the linear system, which reduces the number of iterations by a factor of about three compared to the unpreconditioned case. The number of points per level and the corresponding fill-distances are provided in Table 8.
$\ell$ $N_{\ell}$ $h$ (nested) $h$ (not-nested)
@@
The error is evaluated on a finer set of 2 000 000 points, which are randomly selected once from the original dataset. The evaluation set remains fixed throughout all experiments to ensure consistency in error measurement. The results using the Matérn-1/2 and Matérn-3/2 RBFs are shown in Tables 9 and 10.
@@
Table 9. Results for the interpolation of a test function on the Lucy point cloud for the $C^{0}$-smooth Matérn-1/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
@@
Table 10. Results for the interpolation of a test function on the Lucy point cloud for the $C^{1}$-smooth Matérn-3/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
We observe that the $C^{0}$-smooth Matérn-1/2 RBF performs better in this scenario, primarily due to the improved conditioning of the associated generalized Vandermonde matrix. In both cases, the approximation error reaches the compression error of $10^{-4}$.
A visualization of the solution on all five levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF and not-nested points is depicted in Figure 4.

### v2

@@ 7.1. Franke’s function in two spatial dimensions
on the unit square $\Omega=[0,1]^{2}$. To this end, we generate 10 sets of nested points within this domain, distributing them on a regular grid with size $h_{\ell}=2^{-\ell}$. The number of points per level and the corresponding fill-distances are provided in Table 1.
@@
The solution is approximated with the ${C}^{1}$-smooth Matérn-3/2 RBF and the ${C}^{2}$-smooth Matérn-5/2 RBF for different values of $\gamma$ and using $\rho=2$, $q=4$. Furthermore, we perform an a-posteriori thresholding of small matrix entries with a relative threshold of $10^{-8}$ with respect to the Frobenius norm, cp. Theorem 6.1. The error is evaluated at a fine grid with mesh-size $h=2^{-11}$. The accuracy of the conjugate gradient method is set to $10^{-6}$. Increasing the accuracy does not result in smaller errors but increases the number of iterations. Interpolation error, order, and compression rate (measured as the percentage of nonzero entries, %nz), as well as (non-preconditioned) conjugate gradient iterations (CG), are presented in Tables 25.
@@
4 $6.76\cdot 10^{-3}$ 2.23 48 10 $1.15\cdot 10^{-3}$ 3.00 48 34
5 $1.48\cdot 10^{-3}$ 2.19 22 10 $2.51\cdot 10^{-4}$ 2.53 28 39
6 $3.45\cdot 10^{-4}$ 2.10 7.8 10 $6.08\cdot 10^{-5}$ 2.05 10 39
7 $8.30\cdot 10^{-5}$ 2.06 2.3 10 $1.30\cdot 10^{-5}$ 2.22 3.2 39
8 $1.46\cdot 10^{-5}$ 2.51 0.64 10 $1.14\cdot 10^{-6}$ 3.51 0.89 39
9 $3.13\cdot 10^{-6}$ 2.22 0.17 10 $9.98\cdot 10^{-8}$ 3.52 0.23 39
10 $7.85\cdot 10^{-7}$ 2.00 0.04 10 $3.47\cdot 10^{-8}$ 1.52 0.06 39
Table 2. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
@@
2 $4.67\cdot 10^{-2}$ 3.37 40 17 $4.51\cdot 10^{-2}$ 3.43 40 20
3 $7.97\cdot 10^{-3}$ 2.55 43 40 $7.72\cdot 10^{-3}$ 2.55 43 48
4 $5.70\cdot 10^{-4}$ 3.81 48 63 $3.60\cdot 10^{-4}$ 4.43 48 96
5 $1.09\cdot 10^{-4}$ 2.38 35 90 $5.75\cdot 10^{-5}$ 3.31 37 155
6 $2.38\cdot 10^{-5}$ 2.19 14 96 $1.16\cdot 10^{-5}$ 2.31 15 186
7 $4.57\cdot 10^{-6}$ 2.38 4.4 98 $2.06\cdot 10^{-6}$ 2.49 5.0 195
8 $4.74\cdot 10^{-7}$ 3.26 1.2 99 $2.16\cdot 10^{-7}$ 3.25 1.4 196
9 $3.80\cdot 10^{-8}$ 3.64 0.32 98 $3.05\cdot 10^{-8}$ 2.83 0.38 197
10 $1.33\cdot 10^{-8}$ 1.51 0.08 98 $1.41\cdot 10^{-8}$ 1.11 0.09 200
Table 3. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
@@
2 $1.12\cdot 10^{-1}$ 2.28 40 8 $4.51\cdot 10^{-2}$ 1.39 40 16
3 $2.27\cdot 10^{-2}$ 2.14 43 9 $7.77\cdot 10^{-3}$ 2.54 43 32
4 $4.54\cdot 10^{-3}$ 2.32 48 10 $6.92\cdot 10^{-4}$ 3.48 48 42
5 $1.00\cdot 10^{-3}$ 2.18 22 10 $1.49\cdot 10^{-4}$ 2.21 25 48
6 $2.49\cdot 10^{-4}$ 2.01 7.8 10 $3.42\cdot 10^{-5}$ 2.12 8.8 48
7 $6.13\cdot 10^{-5}$ 2.02 2.3 10 $6.88\cdot 10^{-6}$ 2.31 2.7 48
8 $6.06\cdot 10^{-6}$ 3.34 0.64 9 $8.31\cdot 10^{-7}$ 3.05 0.74 48
9 $1.15\cdot 10^{-6}$ 2.40 0.17 9 $5.68\cdot 10^{-8}$ 3.87 0.19 48
10 $3.07\cdot 10^{-7}$ 1.91 0.04 9 $6.86\cdot 10^{-9}$ 3.05 0.05 48
Table 4. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
@@
2 $4.41\cdot 10^{-2}$ 3.44 40 20 $4.69\cdot 10^{-2}$ 4.29 40 25
3 $7.13\cdot 10^{-3}$ 2.63 43 58 $7.47\cdot 10^{-3}$ 2.65 43 83
4 $3.29\cdot 10^{-4}$ 4.43 48 96 $1.63\cdot 10^{-4}$ 2.19 48 171
5 $6.14\cdot 10^{-5}$ 2.42 35 139 $3.96\cdot 10^{-5}$ 2.04 37 301
6 $1.22\cdot 10^{-5}$ 2.33 13 155 $1.34\cdot 10^{-5}$ 1.56 15 369
7 $2.11\cdot 10^{-6}$ 2.53 4.2 157 $2.06\cdot 10^{-6}$ 2.70 4.8 403
8 $2.57\cdot 10^{-7}$ 3.03 1.2 158 $2.39\cdot 10^{-7}$ 3.10 1.4 407
9 $4.06\cdot 10^{-8}$ 2.66 0.31 159 $7.56\cdot 10^{-8}$ 1.66 0.36 436
10 $3.04\cdot 10^{-8}$ 0.42 0.08 162 $7.01\cdot 10^{-8}$ 0.11 0.09 430
Table 5. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
The tables clearly demonstrate the expected exponential decay of the approximation error with increasing level for both error measures and both RBFs under consideration. Furthermore, we see the effect of a larger value of $\gamma$, which corresponds to a larger lengthscale parameter of the RBF, resulting in a larger condition number. For increasing $\gamma$, we require successively more iterations for the conjugate gradient method to achieve the prescribed accuracy, with up to 430 iterations for $\gamma=1$ and the Matérn-5/2 RBF.
To validate the theoretical bound on the computational cost of assembling the linear system stated in Theorem 2.1, Figure 2 shows the single threaded computation times obtained with $\gamma=1$ and the Matérn-3/2 RBF. The dashed lines correspond to the rates $N\log^{\alpha}N$ for $\alpha=1,2,3$. The results clearly confirm that the overall time scales as $N\log^{2}N$, in agreement with the theoretical analysis.
[…]
Figure 2. Comparison of the overall assembly time (blue) with theoretical growth rates $N\log^{\alpha}N$ for $\alpha=1,2,3$.
Finally, Figure 3 shows a visualization of the solution (left) and the residuals (right) at levels $\ell=1,4,7,10$ for $\gamma=0.5$ using the Matérn-3/2 RBF.
@@ 7.2. Non-smooth function
We consider the L-shaped domain $\Omega=\left[-\frac{1}{2},\frac{1}{2}\right]^{2}\setminus\left(0,\frac{1}{2}\right]^{2}$ where we introduce polar coordinates $x=r\cos\phi$, $y=r\sin\phi$, with $r\geq 0$ and $\phi\in[\pi/2,2\pi]$. In this test, we aim at interpolating the function $u(r,\phi)=-r^{\frac{2}{3}}\sin\left(\frac{2\phi-\pi}{3}\right),$ which solves the Laplace equation $\Delta u=0$. We select $h_{\ell}=2^{-\ell}$ as before. Given the reduced smoothness of the function, we choose the ${C}^{0}$-smooth Matérn-1/2 RBF for its approximation. The error is evaluated on a fine grid with mesh size $h=2^{-9}$. The parameters for the samplet matrix compression and the accuracy of the conjugate gradient method are the same as in the previous example. Table 6 provides the number of points per level along with the corresponding fill distances. The results for $\gamma=0.5$ and $\gamma=1$ are summarized in Table 7.
@@
1 $4.37\cdot 10^{-2}$ $5.98\cdot 10^{-2}$ 52 8
2 $1.38\cdot 10^{-2}$ $3.17\cdot 10^{-2}$ 1.66 0.92 49 15
3 $5.33\cdot 10^{-3}$ $2.04\cdot 10^{-2}$ 1.37 0.64 38 22
4 $1.83\cdot 10^{-3}$ $1.29\cdot 10^{-2}$ 1.54 0.66 18 25
5 $5.88\cdot 10^{-4}$ $8.15\cdot 10^{-3}$ 1.64 0.66 6.2 27
6 $1.85\cdot 10^{-4}$ $5.06\cdot 10^{-3}$ 1.67 0.69 1.8 28
7 $5.84\cdot 10^{-5}$ $3.17\cdot 10^{-3}$ 1.66 0.67 0.51 28
8 $1.84\cdot 10^{-5}$ $1.59\cdot 10^{-3}$ 1.66 1.00 0.13 28
@@
1 $3.39\cdot 10^{-2}$ $4.81\cdot 10^{-2}$ 52 9
2 $1.27\cdot 10^{-2}$ $3.12\cdot 10^{-2}$ 1.42 0.62 51 21
3 $4.15\cdot 10^{-3}$ $1.98\cdot 10^{-2}$ 1.61 0.66 48 37
4 $1.29\cdot 10^{-3}$ $1.25\cdot 10^{-2}$ 1.69 0.66 39 54
5 $3.97\cdot 10^{-4}$ $7.90\cdot 10^{-3}$ 1.70 0.66 19 69
6 $1.22\cdot 10^{-4}$ $4.90\cdot 10^{-3}$ 1.70 0.69 6.7 77
7 $3.79\cdot 10^{-5}$ $3.04\cdot 10^{-3}$ 1.69 0.69 1.9 79
8 $1.13\cdot 10^{-5}$ $1.15\cdot 10^{-3}$ 1.75 1.40 0.44 79
Table 7. Results for the interpolation of the solution of the Laplace equation on the L-shaped domain using the ${C}^{0}$-smooth Matérn-1/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
The compression rates are similar to the previous example, also in case of the less smooth Matérn-1/2 RBF. The approximation error and the corresponding order are reduced, however. The number of conjugate gradient iterations stays rather low with a maximum of 28 iterations for $\gamma=0.5$ and of 79 iterations for $\gamma=1$. This can be explained by the slower decay of the Matérn-1/2 kernel’s eigenvalues, resulting in less ill-conditioned generalized Vandermonde matrices.
@@
Figure 4. Solution (left) and corresponding residuals (right) for the interpolation of the solution to the Laplace equation on the L-shaped domain for levels $\ell=2,4,6,8$.
@@ 7.3. Function on a point cloud in three spatial dimensions
As final example, we consider a function defined on a uniform resampling of the Lucy surface model of the Stanford 3D Scanning Repository. To construct the local approximation spaces, we selected random subsamples of the resulting point cloud, ensuring a specific cardinality for each set. We start with an initial set of 20 points, and at each subsequent level, the cardinality is increased by a factor of 16. In this study, we present results for both nested and non-nested point sets. For the nested case, the indices of the points are randomly shuffled, and subsamples are obtained by selecting the first $n$ points, where $n$ represents the desired cardinality. In contrast, for the non-nested case, each subsample is randomly selected from the entire dataset independently, i.e., without enforcing the inclusion of points from previous levels. Unlike previous tests, this experiment is more general, as the points are not arranged on a uniform grid, they are not necessarily nested, and they are not necessarily quasi-uniform. As test function, we interpolate $f({\boldsymbol{x}})=\big{(}\sqrt[4]{\|{\boldsymbol{x}}-{\boldsymbol{x}}_{0}\|_{2}}+10^{-4}\big{)}^{-1}$ The point ${\boldsymbol{x}}_{0}$ is selected near the torch that Lucy holds in her left hand. To optimize performance, we use a relatively small value for the parameter $\gamma$ and set $\rho=3$, $q=2$, while the a-posteriori threshold is set to $10^{-5}$. These adjustments help decrease both computational time and the number of non-zero entries. However, a very small $\gamma$ can compromise the precision, while a very large $\gamma$ increases the number of conjugate gradient iterations, thereby increasing computational time. A good trade-off is given by $\gamma=0.2$. Furthermore, we employed a diagonally preconditioned conjugate gradient method with accuracy $10^{-6}$ to solve the linear system, which reduces the number of iterations by a factor of about three compared to the unpreconditioned case. The number of points per level and the corresponding fill-distances are provided in Table 8.
$\ell$ $N_{\ell}$ $h$ (nested) $h$ (non-nested)
@@
The error is evaluated on a finer set of 2 000 000 points, which are randomly selected once from the original dataset. The evaluation set remains fixed throughout all experiments to ensure consistency in error measurement. The results using the Matérn-1/2 and Matérn-3/2 RBFs are shown in Tables 9 and 10, respectively.
@@
Table 9. Results for the interpolation of a test function on the Lucy point cloud for the $C^{0}$-smooth Matérn-1/2 RBF using $\rho=3$, $q=2$ as samplet compression parameters.
@@
Table 10. Results for the interpolation of a test function on the Lucy point cloud for the $C^{1}$-smooth Matérn-3/2 RBF using $\rho=3$, $q=2$ as samplet compression parameters.
We observe that the $C^{0}$-smooth Matérn-1/2 RBF performs better in this scenario, primarily due to the improved conditioning of the associated generalized Vandermonde matrix. In both cases, the approximation error is higher than the compression error $\kappa=10^{-5}$.
A visualization of the solution on all five levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF and non-nested points is depicted in Figure 5.